OpenStudy (anonymous):

For the given functions f, g, and h, find f o g o h and state the exact domain of . f(x)=1/x, g(x)=lnx, h(x)=2x+15

OpenStudy (anonymous):

\[h(x)=2x+5\], domain is all real numbers but \[g(x)=\ln(x)\] so domain is \[x>0\] now \[g\circ h(x)=g(h(x))=g(2x+5)=\ln(2x+5)\] so now for domain we have \[2x+5>0\] solving gives \[2x>-5,x>-\frac{5}{2}\] for our new domain

OpenStudy (anonymous):

finally \[f(x)=\frac{1}{x}\] and domain is all numbers except 0, so \[f\circ g\circ h(x)=f\circ \ln(2x+5)=\frac{1}{\ln(2x+5)}\] and now \[\ln(2x+5)\] cannot be zero. that means \[2x+5\neq 1\] because \[\ln(1)=0\] and so we have to solve \[2x+5=1\] \[x=-2\] so domain is \[-\frac{5}{2}<x, x\neq -2\]